OpenStudy (anonymous):

in a party, each member invited takes a number to be served. if the first member took the number 32 and he last member took number 520. How many members were there that day?

OpenStudy (anonymous):

489

OpenStudy (anonymous):

Yep\[\textrm{Last #}-\textrm{First #}+1 = 520-32+1=489\]

OpenStudy (anonymous):

wb and j can complete a lap of jogging track in 6 and 8 minutes respectively. if they start together from the same spot and at the same time then after how many minutes will they reach the starting point together?

OpenStudy (anonymous):

ken is this a rule?

OpenStudy (anonymous):

Post a new question, I'll answer right away

OpenStudy (anonymous):

wb and j can complete a lap of jogging track in 6 and 8 minutes respectively. if they start together from the same spot and at the same time then after how many minutes will they reach the starting point together?

OpenStudy (anonymous):

Oh nvm The answer is LCM(6, 8). Do you know LCM?

OpenStudy (anonymous):

least common factor

OpenStudy (anonymous):

Yea, so what is LCM(6, 8)?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

Yea, so it'll take 24 minutes for them to catch up, 4 laps for wb and 3 for j.

OpenStudy (anonymous):

two sets of 4 consective positive integers have exactly one integer in common. the sum of integers in the set wih greater numbers is how much greater than the sum of integer in the other set?

OpenStudy (anonymous):

ken r u there?

OpenStudy (anonymous):

a number is thrift if it is multiple of 2 or 3. how many thrift numbers are there between -15 and 15?

OpenStudy (anonymous):

If two sets of four consecutive integers have one integer in common, the total in the combined set is 7 and we can write the sets as n + (n + 1) + (n + 2) + (n + 3 ) and (n + 3) + (n + 4) + (n + 5) + (n + 6) Each term in the second set is 3 more than the equivalent term in the first set. Since there are four terms the total of the differences will be 4 x 3 = 12.

OpenStudy (anonymous):

Each term in the second set is 3 more than the equivalent term in the first set. Since there are four terms the total of the differences will be 4 x 3 = 12. i can not understand thi sentence

OpenStudy (anonymous):

Correct about the consecutive-integer sets.\[ \left\{\color{green}{-15,-14,\color{red}{-13}, -12, \color{red}{-11}, -10, -9, -8, \color{red}{-7}, -6, \color{red}{-5}, -4, -3, -2, \color{red}{-1,~0,~1},\\ 2,~3,~4,~\color{red}{5},~6,~\color{red}{7},~8,~9,~10,~\color{red}{11},~12,~\color{red}{13},~14,~15}\right\} \] 20 total thrifts.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

i cant understand plz explain

OpenStudy (anonymous):

19 hn thrift number

OpenStudy (anonymous):

thank you